Grade 12Mechanics

The blades of an aeroplane propeller are 2m long and rotate at 300rpm. Calculate :1. frequency
  1. period of rotation
  2. angular velocity
  3. linear velocity of a point on the blades 0.5m from the tip of the blades.

Aoishi
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
Dear Student,
From the conservation of angular momenta we know,
So at tip of the blade,
I1 *w1 = I2 *w2
where I1 and I2 are the respective angular momentums and w1 and w2 are the angular velicties respectively.
ml^2/3 * 300 = m(0.25l)^2/3 * w
get the value of w from here.\
For Time period it would be: T = 2pi/w
for frequency, f = 1/T
and linear velocity v = w*x where x = 1.5 m from hinge.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)